Name the vertex and axis of symmetry for each quadratic function. Tell wherther the parabola opens up, down, left, or right.
H(x)=-(x-6)^2+1

Answer by RAY100(1637)

Let H(x) = y
,
y = (x-6)^2 +1
,
(y-1) = (x-6)^2,,,,,from this we can read, vertex at (6,1),
,
positive leading coefficient means it is pointing up
,
Axis of symmetry is x=6
,
another form might be,,, y=x^2 -12x +36 +1 = x^2 -12x +37
,
vertex of x is -b/2a = -(-12)/2(1) = +6,,, subst in to get y= 1,,,,( 6,1)
,
check by plotting a few points, (6,1),,,,(5,2),,,,(7,2)